# Quantum optical coherence theory based on Feynman’s path integral

### v1

@@ III.2 First-order interference of two independent light beams
The first-order interference of one light beam confirms Dirac’s statement that photon only interferes with itself. However, based on Popper’s philosophy that a scientific conclusion can not be proved to be true Popper (2005). The first-order interference of light does not prove Dirac’s statement is correct. On the contrary, any scientific conclusion should be possible to be disproved. Dirac’s statement can be disproved if one can observe the first-order interference pattern with two independent light beams.
@@
When Young first discussed interference of light, he pointed out that there is first-order interference pattern of two light beams from different origins by analogy of sound and water waves Young (1802): When two Undulations, from different Origins, coincide either perfectly or very nearly in Direction, their joint effect is a Combination of the Motions belonging to each. Later, Young realized that it is impossible to observe first-order interference pattern of light emitted by two different origins. He observed the first-order interference pattern of sunlight by splitting one beam into two and then combining these two beams together Young (1804). It is now well-accepted that in order to observe the first-order interference pattern with two light beams, these two light beams must have the polarization and frequency, and fixed relative phase Hecht (2012). The easiest way to meet these condition is these two light beams are splitting from the same one. Most of the first-order interferometers are designed in this way Born and Wolf (2013).
@@
Thing becomes different when laser was invented Maiman et al. (1960); Hecht (2010). Several groups superposed two light beams emitted by two independent lasers and observed the first-order interference pattern within the coherence time Javan et al. (1962); Magyar and Mandel (1963); Lipsett and Mandel (1963), which is usually called transient first-order interference pattern. Does the observed transient first-order interference pattern with two independent laser light beams mean that the second part of Dirac’s statement is disproved? Further experiments shows that the first-order interference pattern exists when the intensities of these two laser light beams are so low that there is only one photon in the interferometer in most time Pfleegor and Mandel (1968); Radloff (1968); Hariharan et al. (1993). It is thought by Pfleegor and Mandel that the observed transient first-order interference pattern does not conflict with Dirac’s statement Pfleegor and Mandel (1967). It was the detection of the photon makes the photon into a superposition state and give interference Pfleegor and Mandel (1967). What Glauber said about these experiments is easier to understand Glauber (1995): The things that interfere in quantum mechanics are not particles. They are probability amplitudes for certain events. It is the fact that probability amplitudes add up like a complex numbers that is responsible for all quantum mechanics interferences.
@@ III.2.1 First-order interference of two independent laser light beams
In the first-order interference of two independent laser light beams shown in Fig. 5, there are two different paths for a single-photon detector D detecting a photon. The first one is that the detected photon comes from laser S1 and the corresponding probability amplitude is $A_{1}(x)$. The second one is that the detected photon comes from laser S2 and the corresponding probability amplitude is $A_{2}(x)$. If these two laser light beams have the same polarization, frequency bandwidth, intensity, these two different ways can only be distinguished by momentum difference. As introduced in Sect. II.2.2, when the distance between two lasers satisfies
@@
these two different paths are indistinguishable, where $L$ is the distance between source and observation planes, $\lambda$ is the wavelength of photon, and $\Delta x$ is the position uncertainty of photon detection. The probability function for the $j$th detected photon is
$P_{j}(x)$ $=$ $|A_{j1}(x)+A_{j2}(x)|^{2}$ (30)
$=$ $|e^{i\varphi_{j1}}K_{j1}(x)+e^{i\varphi_{j2}}K_{j2}(x)|^{2},$
@@
Figure 5: The first-order interference of two independent light beams. S1 and S2 are two independent light sources. The distance between two sources is $d$. The distance between the source plane and the observation plane, O, equals $L$. D is a single-photon detector.
For a point light source, Eq. (30) can be simplified as
@@
where paraxial approximation and one-dimension are employed to simplify the calculation. The initial phases of photons in single-mode continuous-wave laser light are identical within the coherence volume. The initial phases of photons in different coherence volumes are random. The initial phases of two independent lasers, $\varphi_{j1}$ and $\varphi_{j2}$, are independent. However, the relative phase, $\varphi_{j1}-\varphi_{j2}$, is fixed within one coherence volume for two independent lasers. The observed intensity is proportional to the sum of the probabilities of all the detected photons
@@
where the first-order interference pattern can not be observed for a long observation time.
The results above are consistent with Magyar and Mandel’s experiments Magyar and Mandel (1963). Even through the physics of the interference of two independent lasers is in question for a long time Mandel (1964); Louradour et al. (1993); Wallace (1994); Davis and Parigger (1994); Paul (1986), the experimental results are certain and the predictions in quantum optical coherence theory is consistent with the one obtained in classical theory.
@@ III.2.2 First-order interference of two independent thermal light beams
Things become interesting for the first-order interference of two independent thermal light beams. As mentioned in Sect. II.3, there are two major differences between photons in laser and thermal light. The first one is that the initial phases of photons in thermal light are random within the coherence volume and the initial phases of photons in laser light are identical within one coherence volume. The second difference is that the degeneracy factor of thermal light is usually much less than 1 and the degeneracy factor of laser light can be much larger than one. The latter difference indicates that it may be impossible to detect large enough number of photons to see whether there is transient first-order interference pattern of two independent thermal light beams within the coherence time. No transient first-order interference pattern of two independent thermal light beams has been reported as the one for two independent lasers Magyar and Mandel (1963).
Theoretically, we can assume the degeneracy factor of thermal light is much larger than one and discuss whether there exists transient first-order interference pattern or not. The scheme for the first-order interference of two independent thermal light beams is also represented by Fig. 5, where S1 and S2 are two identical and independent thermal light sources. There are two different paths for the single-photon detector D detecting a photon. With the same process as the one in Sect. III.2.1, the probability function for the $j$th detected photon is
@@
where paraxial approximation and one-dimension are assumed. The probability function for $N$ detected photon is
@@
where $\varphi$ is the result of the sum of large random phases. The visibility of the first-order interference pattern after detecting $N$ photon is
@@
which decreases as the number of detected photons increases. Figure 6 shows the relationship between the simulated visibility of first-order interference pattern of two independent thermal light beams and the number of detected photons. $V$ is the visibility and $N$ is the number of detected photons. The red line is theoretical value given by Eq. (38) and square dots are numerical results. Theoretical and numerical results are consistent Bai et al. (2016).
Figure 6: The visibility of first-order interference pattern of two independent thermal light beams vs. the number of detected photons Bai et al. (2016).
The first-order interference pattern can not be observed with only one or several photons due to one photon only triggers a photo-electron Dimitrova and Weis (2008). There must be large enough number of photons to retrieve the first-order interference pattern. However, the visibility decreases as the number of detected photon increases via Eq. (38). The first-order interference pattern of two independent thermal light beams may not be observable, which is similar as the first-order interference pattern of two independent laser light beams can not be observed for a long observation time.
@@ III.2.3 First-order interference of two independent non-classical light beams
Feynman’s path integral can also be employed to calculate the first-order interference of two independent non-classical light beams. Assume S1 and S2 in Fig. 5 are two identical and independent single-photon sources. There are two different paths for single-photon detector D detecting a photon, which are emitted by S1 and S2. The probability amplitudes for these two paths are $A_{1}(x)$ and $A_{2}(x)$, respectively. When the distance between these two sources is less than $\lambda L/\Delta x$, these two difference paths can not be distinguished by momentum difference. However, this condition alone does not guarantee that these two paths are indistinguishable as the one in laser and thermal light cases.
@@
$P_{j}(x)$ $=$ $|A_{j1}(x)|^{2}+|A_{j2}(x)|^{2}$ (39)
$=$ $|e^{i\varphi_{j1}}K_{j1}(x)|^{2}+|e^{i\varphi_{j2}}K_{j2}(x)|^{2}.$
The probability equals the sum of the probabilities for different paths. There is no interference.
When these two single-photon sources emit photons simultaneously, there are two photons in the interferometer and these two different paths are distinguishable when all the other parameters are indistinguishable. The probability function for the $j$th detected photon is
$P_{j}^{\prime}{}(x)=|e^{i\varphi_{j1}}K_{j1}(x)+e^{i\varphi_{j2}}K_{j2}(x)|^{2}.$ (40)
The probability function after detecting $N$ photons is the sum of the probabilities correspond to these two cases,
@@
$N_{2}$ is much less than $N$ for two independent single-photon sources. The visibility of the first-order interference pattern of two independent single-photon states is much less than the one of two-independent thermal light beams. It is impossible to observe the first-order interference pattern of two independent single-photon states for either short or long time comparing to the coherence time.
With the same method above, one can calculate the first-order interference of two independent number states with different numbers, one thermal and one laser light beams, and other conditions. No transient first-order interference pattern can be observed except the one of two independent laser light beams. No first-order interference pattern can be observed for two independent light beams with long observing time.
III.2.4 Remakes about the classical model of laser and thermal light field
With the results above, it is interesting to re-visit the electric model for laser and thermal light in classical electromagnetic theory. For simplicity, quasi-monochromatic light and plane wave are assumed. The electric field for single-mode continuous-wave laser light at space-time coordinate $(\vec{r},t)$ is Loudon (2000); Born and Wolf (2013)
@@ III.2.4 Remakes about the classical model of laser and thermal light field
The same model as Eq. (43) is employed to describe thermal light field within one coherence time. For instance, in Born and Wolf’s book, they wrote: In a monochromatic wave field the amplitude of the vibrations at any point P is constant, while the phase varies linearly with time. This is no longer the case in a wave field produced by a real source: the amplitude and phase undergo irregular fluctuation, the rapidity of which depends essentially on the effective width $\Delta\nu$ of the spectrum. The complex amplitude remains substantially constant only during a time interval $\Delta t$ which is small compared to the reciprocal of the effective spectral width $\Delta\nu$; in such a time interval the change of the relative phase of any two Fourier components is much less than $2\pi$ and the addition of such components represents a disturbance which in this time interval behaves like a monochromatic wave with the mean frequency. However, this is not true for a longer time interval. The characteristic time $\Delta t=1/\Delta v$ is of the order of the coherence time introduced in §7.5.8. They thought the amplitude and phase of electric field produced by a real source undergoes irregular fluctuations for a long time interval. The amplitude and phase of electric field can be seen as constant within the coherence time. In other words, the electric field produced by real source, i.e., laser or thermal light sources, can be described by Eq. (43). The same electric field model is employed to describe laser and thermal light field within the coherence time. This model is employed by Forrester to calculate the first-order interference of two independent thermal light beams, too Forrester (1956).
@@
In classical theory, it is straightforward to obtain the conclusion that there is transient first-order interference pattern by superposing two independent thermal light beams as the one of two independent laser light beams. This conclusion is in conflict with the conclusion obtained in Feynman’s path integral in Sect. III.2.2. Experiments should be employed to verify the different conclusions. However, due to the very low degeneracy factor of thermal light, it is very difficult to observe transient first-order interference pattern of two independent thermal light beams directly.
@@
Forrester et al. claimed they observed interference between independently generated light waves Forrester et al. (1955). Detail analysis of their experiments indicates that their experiments is a type of one beam interference including nonlinear effect as the slit in a Young’s double-slit interferometer, which is similar as the ones with single atom Grangier et al. (1985); Eichmann et al. (1993) and atom ensembles Afzelius et al. (2007) as the slit in a Young’s double-slit interferometer. Forrester et al.’s experiments can not be interpreted as the first-order interference of two independent thermal light beams.
@@
If there is no transient first-order interference pattern with two independent thermal light beams, the classical model for thermal light within the coherence time may not be the same as the one of laser light. This result is helpful for us to understand the physics of thermal light and worthy of further experimental studies.
@@ III.3 First-order interference of multiple independent light beams
The same method can be employed to calculate the first-order interference of multiple independent light beams. Figure 7 shows the first-order interference of three light beams emitted by three independent light sources, S1, S2, and S3.
Figure 7: First-order interference of three independent light beams. S1, S2, and S3 are three independent light sources. D is a single-photon detector in the observation plane. $L$ is the distance between the source and observation planes.
There are three different paths for the single-photon detector D detecting a photon. The first one is that the detected photon is emitted by S1 and the corresponding probability amplitude is $A_{1}(S,D))$. The second and third paths can be defined similarly and the corresponding probability amplitudes are $A_{2}(S,D)$ and $A_{3}(S,D)$, respectively. For simplicity, these three light sources are assumed to be identical single-mode continuous-wave lasers with equal intensities. The polarizations and frequency distributions of the photons emitted by these three lasers are assumed to be the same. When the distance between $S_{1}$ and S3 is less than $\lambda L/\Delta x$, these three different paths to trigger a photon detection event are indistinguishable. The probability function for the $j$th detected photon is
@@
$=$ $|A_{1}(S,D)+A_{2}(S,D)+A_{3}(S,D)|^{2}$
$=$ $|e^{i\varphi_{j1}}K_{1}(S,D)+e^{i\varphi_{j2}}K_{2}(S,D)+e^{i\varphi_{j3}}K_{3}(S,D)|^{2},$
where $\varphi_{j1}$, $\varphi_{j2}$ and $\varphi_{j3}$ are the initial phases of the $j$th detected photon emitted by S1, S2 and S3, respectively. $K_{1}(S,D)$, $K_{2}(S,D)$ and $K_{3}(S,D)$ are the corresponding photon’s Feyman propagators for these three paths.
Equation (7) can be simplified as
@@
where paraxial approximation and one-dimension are assumed to simplify the calculation, $d_{12}$ is the distance between S1 and S2, similar definitions hold for $d_{13}$ and $d_{23}$. The relative position for these three sources is shown in Fig. 8.
Figure 8: Relative positions of three sources.
Since these three light sources are independent lasers, $\varphi_{j1}$, $\varphi_{j2}$ and $\varphi_{j3}$ are random. The relative phases are fixed if all the detected photons are within one coherence volume. The probability for $N$ detected photon is
@@
where transient first-order interference pattern between three independent laser light beams can be observed. Figure 9 shows transient first-order interference pattern of three laser light beams with equal initial phases. When $d_{12}\neq d_{23}$, the visibility of the transient first-order interference pattern can not reach 100%. The visibility can reach 100% when $d_{12}=d_{23}$ as shown in Fig. 9(b).
Figure 9: Transient first-order interference pattern for (a) $d_{12}\neq d_{23}$ and (b) $d_{12}=d_{23}$.
When the observation time is much longer than the coherence time, the probability for $N$ detected photons can be approximated as a constant due to the random initial phases.
The same method can be employed to calculate the first-order interference of three and more different types of independent light beams. It can be concluded that only when more than two laser light beams are employed, transient first-order interference pattern can be observed.

### v2

@@ III.2 First-order interference of two independent light beams
The first-order interference of one light beam confirms Dirac’s statement that photon only interferes with itself. Based on Popper’s philosophy that a scientific conclusion can not be proved to be true Popper (2005). The first-order interference of one light beam does not prove Dirac’s statement is correct. On the contrary, any scientific conclusion should be possible to be disproved. Dirac’s statement may be disproved if one can observe the first-order interference pattern with two independent light beams.
@@
When Young first discussed the interference of light, he pointed out that there is first-order interference pattern of two light beams from different origins by analogy of sound waves Young (1802): When two Undulations, from different Origins, coincide either perfectly or very nearly in Direction, their joint effect is a Combination of the Motions belonging to each. Later, Young realized that it is impossible to observe first-order interference pattern of light from two different origins. He observed the first-order interference pattern of sunlight by splitting one beam into two and then combining these two beams together Young (1804). It is now well-accepted that in order to observe the first-order interference pattern with two light beams, these two light beams must have the same polarization and frequency, and fixed relative phase Hecht (2012). The easiest way to meet these conditions is that these two light beams are splitting from the same one. Most of the first-order interferometers are designed in this way Born and Wolf (2013).
@@
Thing becomes different when laser was invented Maiman et al. (1960); Hecht (2010). Several groups superposed two light beams emitted by two independent lasers and observed the first-order interference pattern within the coherence time Javan et al. (1962); Magyar and Mandel (1963); Lipsett and Mandel (1963), which is usually called transient first-order interference pattern. Does the observed transient first-order interference pattern with two independent laser light beams means that the second part of Dirac’s statement is disproved? Further experiments shows that the first-order interference pattern exists when the intensities of these two laser light beams are so low that there is only one photon in the interferometer in most time Pfleegor and Mandel (1968); Radloff (1968); Hariharan et al. (1993). It is interpreted by Pfleegor and Mandel that the observed transient first-order interference pattern does not conflict with Dirac’s statement Pfleegor and Mandel (1967). It is the detection of a photon that puts the photon into a superposition state and causes interference Pfleegor and Mandel (1967). What Glauber said about these experiments is easier to understand Glauber (1995): The things that interfere in quantum mechanics are not particles. They are probability amplitudes for certain events. It is the fact that probability amplitudes add up like a complex numbers that is responsible for all quantum mechanics interferences.
@@ III.2.1 First-order interference of two independent laser light beams
In the scheme of the first-order interference of two independent laser light beams shown in Fig. 5, there are two different paths for the single-photon detector, D, detecting a photon. The first path is that the detected photon comes from laser S1. The probability for the detected photon comes from S1 is $I_{1}/(I_{1}+I_{2})$, where $I_{1}$ and $I_{2}$ are the intensities of light emitted by S1 and S2, respectively. The corresponding probability amplitude is $A_{1}(x)$. The second one is that the detected photon comes from laser S2. The probability for the detected photon comes from S2 is $I_{2}/(I_{1}+I_{2})$ and the corresponding probability amplitude is $A_{2}(x)$. If these two laser light beams have the same polarization and frequency bandwidth, these two different paths can only be distinguished by momentum difference. As introduced in Sect. II.2.2, when the distance between two lasers satisfies
@@
these two different paths are indistinguishable, where $L$ is the distance between source and observation planes, $\lambda$ is the wavelength of the photon, and $\Delta x$ is the position uncertainty of photon detection. The probability function of the $j$th detected photon is
$P_{j}(x)$ $=$ $|\sqrt{\frac{I_{1}}{I_{1}+I_{2}}}A_{j1}(x)+\sqrt{\frac{I_{2}}{I_{1}+I_{2}}}A_{j2}(x)|^{2}$ (30)
$\propto$ $|\sqrt{I_{1}}e^{i\varphi_{j1}}K_{j1}(x)+\sqrt{I_{2}}e^{i\varphi_{j2}}K_{j2}(x)|^{2},$
@@
Figure 5: The first-order interference of two independent light beams. S1 and S2 are two independent light sources. The distance between two sources is $d$. The distance between the source and observation planes equals $L$. O is the observation plane. D is a single-photon detector.
For two point light sources with equal intensity, Eq. (30) can be simplified as
@@
where paraxial approximation and one-dimension are employed to simplify the calculation Born and Wolf (2013). The initial phases of photons in a single-mode continuous-wave laser light beam are identical within one coherence volume. The initial phases of photons in different coherence volumes are random. The initial phases of two independent lasers, $\varphi_{j1}$ and $\varphi_{j2}$, are independent. However, the relative phase, $\varphi_{j1}-\varphi_{j2}$, is fixed within one coherence volume for two independent lasers. The observed intensity is proportional to the sum of the probability functions of all the detected photons
@@
where the first-order interference pattern of two independent laser light beams disappears for a long observation time.
The results above are consistent with Magyar and Mandel’s experiments Magyar and Mandel (1963). Even through the physics of the interference of two independent lasers is in question for a long time Mandel (1964); Louradour et al. (1993); Wallace (1994); Davis and Parigger (1994); Paul (1986), the experimental results are certain and the predictions in quantum optical coherence theory is consistent with the ones obtained in classical optical coherence theory.
@@ III.2.2 First-order interference of two independent thermal light beams
Things become different for the first-order interference of two independent thermal light beams. There are two major differences between photons in laser and thermal light. The first difference is that the initial phases of photons in thermal light are random within one coherence volume and the initial phases of photons in laser light are identical within one coherence volume. The second difference is that the degeneracy factor of thermal light is usually much less than one and the degeneracy factor of laser light can be much larger than one. The latter difference indicates that it may be impossible to detect large enough number of photons to observe whether there is transient first-order interference pattern of two independent thermal light beams within the coherence time or not. No experimental observation of the transient first-order interference pattern of two independent thermal light beams was reported as the one for two independent lasers Magyar and Mandel (1963).
Theoretically, one can assume that the degeneracy factor of thermal light is much larger than one and discuss whether there exists transient first-order interference pattern or not. The scheme for the first-order interference of two independent thermal light beams is shown in Fig. 5, where S1 and S2 are two identical and independent thermal light sources. There are two different paths for the single-photon detector D detecting a photon. Assume the intensities of these two light beams are equal and the same assumption is always employed in the following situations unless specified otherwise. With the same process as the one in Sect. III.2.1, the probability function for the $j$th detected photon is
@@
where paraxial approximation and one-dimension are assumed. The probability function for $N$ detected photons is
@@
where $\varphi$ is the result of the sum of large number of random phases. The visibility of the first-order interference pattern after detecting $N$ photon is
@@
which decreases as the number of detected photons increases. Figure 6 shows the relationship between the visibility of the simulated first-order interference pattern of two independent thermal light beams and the number of detected photons. $V$ is the visibility and $N$ is the number of detected photons. The red solid line is theoretical value given by Eq. (38) and square dots are numerical results. Theoretical and numerical results are consistent Bai et al. (2016).
Figure 6: The visibility of first-order interference pattern of two independent thermal light beams vs. the number of detected photons Bai et al. (2016). $V$ is the visibility of first-order interference pattern and $N$ is the number of detected photons.
The first-order interference pattern of light can not be observed with only one or several photons due to one photon only triggers one photo-electron Dimitrova and Weis (2008). There must be large enough number of photons to retrieve the first-order interference pattern. However, the visibility decreases as the number of detected photons increases via Eq. (38). The first-order interference pattern of two independent thermal light beams may not be observable. The reason is similar as the one for the first-order interference pattern of two independent laser light beams can not be observed for a long observation time.
@@ III.2.3 First-order interference of two independent non-classical light beams
Feynman’s path integral can also be employed to calculate the first-order interference of two independent non-classical light beams. Assume S1 and S2 in Fig. 5 are two identical and independent single-photon sources. There are two different paths for single-photon detector detecting a photon. The detected photon is emitted by either S1 or S2. The probability amplitudes for these two paths are $A_{1}(x)$ and $A_{2}(x)$, respectively. When the distance between these two sources is less than $\lambda L/\Delta x$, these two difference paths can not be distinguished by momentum difference. However, these conditions do not guarantee that these two paths are indistinguishable as the one in laser and thermal light cases.
@@
$P_{j}(x)$ $=$ $|\frac{1}{\sqrt{2}}A_{j1}(x)|^{2}+|\frac{1}{\sqrt{2}}A_{j2}(x)|^{2}$ (39)
$\propto$ $|e^{i\varphi_{j1}}K_{j1}(x)|^{2}+|e^{i\varphi_{j2}}K_{j2}(x)|^{2}.$
The probability function equals the sum of the probability functions for these two paths. There is no interference.
When these two single-photon sources emit photons simultaneously, there are two photons in the interferometer and these two different paths can be indistinguishable when all the other conditions are satisfied. The probability function for the $j$th detected photon is
$P_{j}^{\prime}{}(x)=|\frac{1}{\sqrt{2}}e^{i\varphi_{j1}}K_{j1}(x)+\frac{1}{\sqrt{2}}e^{i\varphi_{j2}}K_{j2}(x)|^{2}.$ (40)
The probability function after detecting $N$ photons is the sum of the probability function corresponding to these two cases,
@@
$N_{2}$ is much less than $N$ for the probability of two independent single-photon sources simultaneously emitting one photon each is much less than the probability of these two sources emitting one photon only. The visibility of the first-order interference pattern of two independent single-photon states is much less than the one of two-independent thermal light beams. It is impossible to observe the first-order interference pattern of two independent single-photon states for either short or long time comparing to the coherence time.
With the same method above, one can calculate the first-order interference of two independent number states with different numbers of photons, one thermal and one laser light beams, and so on. No transient first-order interference pattern can be observed except the one of two independent laser light beams. No first-order interference pattern exists for two independent light beams with long observing time.
III.2.4 Remakes about the classical models of laser and thermal light field
With the results above, it is interesting to re-visit the electric field models for laser and thermal light in classical electromagnetic theory. For simplicity, quasi-monochromatic light and plane wave are assumed. The electric field for single-mode continuous-wave laser light at space-time coordinate $(\vec{r},t)$ is Loudon (2000); Born and Wolf (2013)
@@ III.2.4 Remakes about the classical models of laser and thermal light field
The same model as Eq. (43) is employed to describe electric field of thermal light within one coherence time. For instance, in Born and Wolf’s book Born and Wolf (2013), they wrote: In a monochromatic wave field the amplitude of the vibrations at any point P is constant, while the phase varies linearly with time. This is no longer the case in a wave field produced by a real source: the amplitude and phase undergo irregular fluctuation, the rapidity of which depends essentially on the effective width $\Delta\nu$ of the spectrum. The complex amplitude remains substantially constant only during a time interval $\Delta t$ which is small compared to the reciprocal of the effective spectral width $\Delta\nu$; in such a time interval the change of the relative phase of any two Fourier components is much less than $2\pi$ and the addition of such components represents a disturbance which in this time interval behaves like a monochromatic wave with the mean frequency. However, this is not true for a longer time interval. The characteristic time $\Delta t=1/\Delta v$ is of the order of the coherence time introduced in §7.5.8. It means that the amplitude and phase of electric field produced by a real source undergo irregular fluctuations for a long time interval. The amplitude and phase of electric field can be seen as constant within one coherence time. In other words, the electric field produced by real source, i.e., laser or thermal light sources, can be described by Eq. (43). The same electric field model is employed to describe laser and thermal light field within the coherence time. This model was employed by Forrester to calculate the first-order interference of two independent thermal light beams, too Forrester (1956).
@@
In classical theory, it is straightforward to obtain the conclusion that there is transient first-order interference pattern by superposing two independent thermal light beams as the one of two independent laser light beams. This conclusion is in conflict with the conclusion obtained in Feynman’s path integral in Sect. III.2.2. Experiments should be employed to verify these two different conclusions. However, due to the very low degeneracy factor of thermal light, it is difficult to observe transient first-order interference pattern of two independent thermal light beams directly.
@@
Forrester et al. claimed that they observed interference between independently generated light waves Forrester et al. (1955). Detail analysis of their experiments indicates that their experiment belongs to one beam interference including nonlinear effect as the slits in a Young’s double-slit interferometer, which is similar as the ones with single atom Grangier et al. (1985); Eichmann et al. (1993) and atom ensembles Afzelius et al. (2007) as the slits in a Young’s double-slit interferometer. Forrester et al.’s experiments can not be interpreted as the results of first-order interference of two independent thermal light beams.
@@
If there is no transient first-order interference pattern with two independent thermal light beams, the classical model for electric field of thermal light within the coherence time may not be the same as the one of laser light. This result is helpful to understand the physics of thermal light and be worthy of further experimental studies.
@@ III.3 First-order interference of multiple independent light beams
The same method can be employed to calculate the first-order interference of multiple independent light beams. Figure 7 shows the scheme for the first-order interference of three light beams emitted by three independent light sources, S1, S2, and S3.
Figure 7: First-order interference of three independent light beams. S1, S2, and S3 are three independent light sources. D is a single-photon detector in the observation plane, O. $L$ is the distance between the source and observation planes.
There are three different paths for the single-photon detector, D, detecting a photon. The first path is that the detected photon is emitted by S1 and the corresponding probability amplitude is $A_{1}(S,D)$. The second and third paths are defined similarly and the corresponding probability amplitudes are $A_{2}(S,D)$ and $A_{3}(S,D)$, respectively. For simplicity, these three light sources are assumed to be identical single-mode continuous-wave lasers with equal intensities. The polarizations and frequency bandwidths of the photons emitted by these three lasers are the same. When the distance between $S_{1}$ and S3 is less than $\lambda L/\Delta x$, these three different paths to trigger a photon detection event are indistinguishable. The probability function for the $j$th detected photon is
@@
$=$ $|\frac{1}{\sqrt{3}}[A_{1}(S,D)+A_{2}(S,D)+A_{3}(S,D)]|^{2}$
$\propto$ $|e^{i\varphi_{j1}}K_{1}(S,D)+e^{i\varphi_{j2}}K_{2}(S,D)+e^{i\varphi_{j3}}K_{3}(S,D)|^{2},$
where $\varphi_{j1}$, $\varphi_{j2}$ and $\varphi_{j3}$ are the initial phases of the $j$th detected photon emitted by S1, S2 and S3, respectively. $K_{1}(S,D)$, $K_{2}(S,D)$ and $K_{3}(S,D)$ are the corresponding photon’s Feynman propagators for the three paths.
Equation (III.3) can be simplified as
@@
where paraxial approximation and one-dimension case are assumed to simplify the calculation, $d_{12}$ is the distance between S1 and S2, similar definitions hold for $d_{13}$ and $d_{23}$. The relative positions for the three sources are shown in Fig. 8.
Figure 8: Relative positions of three sources S1, S2 and S3.
Since the three lasers are independent, $\varphi_{j1}$, $\varphi_{j2}$ and $\varphi_{j3}$ are random. The relative phases are fixed if all the detected photons are within one coherence volume. The probability function for $N$ detected photons is
@@
where transient first-order interference pattern of three independent laser light beams can be observed. Figure 9 shows the transient first-order interference pattern of three laser light beams with equal initial phases. When $d_{12}\neq d_{23}$, the visibility of the transient first-order interference pattern can not reach 100% as shown in Fig. 9(a). The visibility can reach 100% when $d_{12}=d_{23}$ as shown in Fig. 9(b).
Figure 9: Transient first-order interference patterns (a) $d_{12}\neq d_{23}$ and (b) $d_{12}=d_{23}$.
When the observation time is much longer than the coherence time, the probability function for $N$ detected photons is approximated to be a constant due to the sum of large number of random initial phases.
The same method can be employed to calculate the first-order interference of three and more different types of independent light beams. It can be predicted that only when two or more laser light beams are employed, transient first-order interference pattern can be observed.